Charles Austin of the controller’s office of Ayayai Corporation was given the assignment of determining the basic and diluted earnings per share values for the year ending December 31, 2021. Austin has compiled the information listed below.

1. The company is authorized to issue 8,120,000 shares of $10 par value common stock. As of December 31, 2020, 2,030,000 shares had been issued and were outstanding.
2. The per share market prices of the common stock on selected dates were as follows.

Price per Share

July 1, 2020 $20.00
January 1, 2021 21.00
April 1, 2021 25.00
July 1, 2021 11.00
August 1, 2021 10.50
November 1, 2021 9.00
December 31, 2021 10.00
3. A total of 753,600 shares of an authorized 1,164,000 shares of convertible preferred stock had been issued on July 1, 2020. The stock was issued at its par value of $25.00, and it has a cumulative dividend of $3 per share. The stock is convertible into common stock at the rate of one share of convertible preferred for one share of common. The rate of conversion is to be automatically adjusted for stock splits and stock dividends. Dividends are paid quarterly on September 30, December 31, March 31, and June 30.
4. Ayayai Corporation is subject to a 20% income tax rate.
5. The after-tax net income for the year ended December 31, 2021, was $11,840,000.


The following specific activities took place during 2021.

1. January 1—A 5% common stock dividend was issued. The dividend had been declared on December 1, 2020, to all stockholders of record on December 29, 2020.
2. April 1—A total of 439,200 shares of the $3 convertible preferred stock was converted into common stock. The company issued new common stock and retired the preferred stock. This was the only conversion of the preferred stock during 2021.
3. July 1—A 2-for-1 split of the common stock became effective on this date. The board of directors had authorized the split on June 1.
4. August 1—A total of 298,800 shares of common stock were issued to acquire a factory building.
5. November 1—A total of 25,200 shares of common stock were purchased on the open market at $9.00 per share. These shares were to be held as treasury stock and were still in the treasury as of December 31, 2021.
6. Common stock cash dividends—Cash dividends to common stockholders were declared and paid as follows.
        April 15—$0.30 per share
        October 15—$0.20 per share
7. Preferred stock cash dividends—Cash dividends to preferred stockholders were declared and paid as scheduled.


(a) Determine the number of shares used to compute basic earnings per share for the year ended December 31, 2021. (Round answer to 0 decimal places, e.g. $1,500.)

Number of shares to compute basic earnings per share


(b) Determine the number of shares used to compute diluted earnings per share for the year ended December 31, 2021. (Round answer to 0 decimal places, e.g. $1,500.)

Number of shares to compute diluted earnings per share


(c) Compute the adjusted net income to be used as the numerator in the basic earnings per share calculation for the year ended December 31, 2021.

Adjusted net income $
Answer #1

Part A

Event

Dates outstanding

Shares outstanding

restatement

Fraction of year

Weighted shares

Beginning Balance, including 5% stock dividend

Jan. 1–Apr. 1

2131500

2.0

3/12

1065750

Conversion of preferred stock

Apr. 1–July 1

2592660

2.0

3/12

1296330

Stock split

July 1–Aug. 1

5185320

1/12

432110

Issued shares for building

Aug. 1–Nov. 1

5484120

3/12

1371030

Purchase of Treasury stock

Nov. 1–Dec. 31

5458920

2/12

909820

Total number of common shares to compute basic earnings per share

5075040

2030000*1.05 = 2131500

(2131500+439200+(439200*5%)) = 2592660

2592660*2 = 5185320

5185320+298800 = 5484120

5484120-25200 = 5458920

Part 2

Number of shares to compute basic earnings per

5075040

Convertible preferred stock—still outstanding ((298800) X 2 X 1.05)

627480

Convertible preferred stock—converted (439200 X 2 X 1.05 X 3/12)

230580

Number of shares to compute diluted earnings per share

5933100

Part 3

After-tax net income

11840000

Preferred stock dividends

March 31 (753600 X $0.75)

565200

June 30, September 30, and December 31 (314400 X $0.75 X 3)

707400

(1272600)

Adjusted net income

$10567400

753600-439200 = 314400